A Carnot engine operating between 900℉ and 90℉ produces 40,000 ft.lb of work. Determine (a) the heat supplied, (b) the change of entropy during heat rejection, (c) the thermal efficiency of the engine.


1
Expert's answer
2021-11-24T11:57:47-0500

a)

"Q_2=Q_1\frac{T_2}{T_1}=99000~ft.lb,"

b)

"S=\frac{Q_2}{\Delta T}=298~\frac JK,"

c)

"\eta=1- \frac{T_2}{T_1}=0.5=50\%."